# Attention-Seeker: Dynamic Self-Attention Scoring for Unsupervised Keyphrase Extraction

### v1

@@ 7 Limitations
Attention-Seeker estimates the relevance of attention vectors by defining a set of desired characteristics and measuring how closely each attention vector matches them. However, the current implementation of this approach present two major limitations. First, it uses the dot product between the desired vector and SAMs’ attention vectors. While effective, this approach could be improved by applying normalization techniques such as $l_{2}$-normalization (cosine similarity), softmax normalization (cross-attention), or softmax with temperature scaling. Second, the hypothesized vectors ($H$), which serve as reference vectors, are initially defined as binary vectors that attend to candidate phrases. This simplistic formulation may introduce noise into the attention distribution of the candidates. Although we propose a refinement of this vector ($H$) using the attention scores extracted from the document ($B$), the initial estimation of $B$ still relies on the original binary vector. To address these limitations, future work could explore more sophisticated hypothesized vectors and more effective methods for measuring their alignment with SAM’s attention vectors.
Another important limitation is that the long document version of Attention-Seeker requires the first segment to be the abstract of the document. This condition limits the application of our proposed method to documents with an abstract. In addition, our ablation study showed that this approach negatively affects the overall performance of our method. Accordingly, we believe that further research should explore alternative methods for defining the hypothesized vector $H^{s}$ without relying on information extracted from the abstract. Possible solutions could include extracting the most important tokens from all segments under the assumption of equal relevance, and defining the hypothesized vector $H^{s}$ based on these tokens.
@@ Appendix A Optimal Performance of SAMRank on LLAMA 3-8B
The independent implementation of the Global Score in SAMRank outperforms Attention-Seeker in short documents, suggesting room for improvement in our method. As discussed in the main paper, these improvements could include more effective methods for defining the desired characteristics of attention vectors and better measuring their alignment with these characteristics. While we might expect results similar to or better than those reported in Table  5 (Global Score), we must also consider the possibility that non-relevant SAMs might achieve high performance by chance due to the use of non-ideal labels and a large set of 1024 trials.
@@
Appendix B Results of Non-Parametric SAMRank
@@
We implemented the non-parametric SAMRank method using the model LLAMA 3-8B, evaluating their three different scores to ensure that the non-parametric version benefits from these scores similarly to the original SAMRank. Table  6 shows the performance achieved by the non-parametric SAMRank in the three cases.
@@
Appendix C Most Relevant SAMs identified by Attention-Seeker
@@
Figure  3 shows the distribution of $R^{lh}$ scores for each layer, with the scores normalized for consistent comparison. For better visualization, the top 5% of outliers are filtered out from each distribution, distributions are color-coded according to their relevance rank, and a red line is added at $Relevance=3.125$ to indicate a uniform distribution where all layers are equally relevant. The analysis shows that the heads between layers 8 and 18, especially layers 9, 12, 14, and 15, received the highest relevance scores. This result suggests that heads located in intermediate layers of an LLM tend to specialize in processing tasks related to noun-noun relationships and selecting the most relevant nouns/phrases.
@@
Figure  4 supports the previous observation by illustrating the relevance scores $R^{lh}$ assigned to each head for a sample from the Inspec dataset. The score distribution across layers aligns with the general pattern in Figure  3, where most heads in the intermediate layers contribute significantly to keyphrase extraction, while only a few heads in the first and last layers exhibit high relevance. For the same Inspec sample, we compared the attention scores of candidate phrase tokens in layer 3 (lowest overall relevance score) and layer 14 (highest overall relevance score). Figure  5 shows the difference between these two layers: Layer 14 assigns higher attention scores to all candidates, with a clearer contrast between phrases, while layer 3 tends to focus on non-candidate tokens, resulting in lower, more uniformly distributed attention scores.

### v2

@@ 7 Limitations
Attention-Seeker estimates the relevance of attention vectors by defining a set of desired characteristics and measuring how closely each attention vector matches them. However, the current implementation of this approach presents two major limitations. First, it uses the dot product between the desired vector and SAMs’ attention vectors. While this method is effective, it could be improved by applying normalization techniques such as $l_{2}$-normalization (cosine similarity), softmax normalization (cross-attention), or softmax with temperature scaling.
Second, the hypothesis vectors ($H$), which serve as reference vectors, are initially defined as binary vectors that focus on candidate phrases. This simplistic formulation may introduce noise into the attention distribution of the candidates and partially overlook important contextual information. Although we refine this vector ($H$) by using the attention scores extracted from the document ($B$), the initial estimation of $B$ still relies on the original binary vector. To address these limitations, future work could explore more sophisticated hypothesis vectors and improved methods for measuring their alignment with SAM’s attention vectors.
Another important limitation is that the long document version of Attention-Seeker requires the first segment to be the abstract of the document. This condition limits the application of our proposed method to documents with an abstract. In addition, our ablation study showed that this approach negatively affects the overall performance of our method. Accordingly, we believe that further research should explore alternative methods for defining the hypothesis vector $H^{s}$ without relying on information extracted from the abstract. Possible solutions could include extracting the most important tokens from all segments under the assumption of equal relevance, and defining the hypothesis vector $H^{s}$ based on these tokens.
Finally, the implementation of our method is currently limited to open LLMs. Since Attention-seeker relies on the information provided by the Self-Attention Maps of Language Models, it cannot be applied to private models with restricted access to their weights, such as Open AI’s GPT-4o/Open-o1 and Antropic’s Claude models. This limitation is common in the field, as previous methods also rely on internal model information (e.g., SAMs, embeddings, logits). However, as open models continue to improve and close the performance gap with closed models, we believe our approach will provide valuable insights for future research.
Acknowledgments
This work was supported by JST CREST Grant Number JPMJCR22D1 and JSPS KAKENHI Grant Number JP22H00551, Japan.
@@ Appendix A Optimal Performance of SAMRank on LLAMA 3-8B
The independent implementation of the Global Score in SAMRank outperforms Attention-Seeker in short documents, suggesting room for improvement in our method. As discussed in the main paper, these improvements could include more effective strategies for defining the desired characteristics of attention vectors and better measuring their alignment with these characteristics. In future implementations, we might expect results similar to those reported in Table  5 (Global Score). However, it is important to note that these results may be inflated by non-relevant SAMs achieving high performance by chance, due to suboptimal labels and the large number of 1024 possible trials.
Method Inspec SemEval2017 SemEval2010 Krapivin
F1@5 F1@10 F1@15 F1@5 F1@10 F1@15 F1@5 F1@10 F1@15 F1@5 F1@10 F1@15
BERT Model
SAMRank (B) 32.43 38.26 38.84 22.93 32.14 35.85 15.19 18.70 20.21 11.99 12.51 12.14
Attention-Seeker 34.54 40.31 40.15 24.13 34.17 38.30 17.27 21.53 23.05 19.09 17.60 15.56
SAMRank (O) 33.96 39.35 39.73 24.08 33.40 37.53 15.28 18.36 18.03 16.35 15.91 14.52
GPT-2 Model
SAMRank (B) 33.96 38.24 38.11 23.99 32.51 35.75 17.23 20.76 21.56 13.37 15.19 14.27
Attention-Seeker 34.14 39.46 38.89 24.30 33.83 37.16 18.82 23.08 23.73 19.89 17.95 15.80
SAMRank (O) 33.92 39.44 39.72 24.80 34.75 38.78 15.88 18.36 19.03 17.49 16.46 14.92
Table 6: Performance of keyphrase extraction of SAMRank and Attention-Seeker on Small Language Models. SAMRank (B):non-parametric base method. SAMRank (O): optimal manual selection (Kang and Shin, 2023).
@@
Appendix B Performance in Small Language Models: BERT and GPT-2
We replicated the evaluation of Attention-Seeker and the non-parametric SAMRank method (Base) from the main paper, applying our method to smaller Language Models, specifically BERT (Devlin et al., 2019) and GPT-2 (Radford et al., 2019). The results of this evaluation are summarized in Table 6.
These results align with those obtained using the LLAMA 3-8B implementation (Table 2) and indicate that Attention-Seeker always outperforms the non-parametric SAM (base model that treats all layers and heads as equally relevant; Appendix C). These findings emphasize the importance of considering the relevance of an LLM’s layers and heads in the process of keyphrase extraction.
Similar to the analysis in Appendix A, we further compare the results of Attention-Seeker with the optimal performance of SAMRank reported by Kang and Shin (2023). For short documents, the optimal results from SAMRank suggest room for improvement in our method, while for long documents, Attention-Seeker consistently outperforms SAMRank. Although these results are similar to those obtained with the LLAMA 3-8B implementation, Attention-Seeker outperforms the optimal SAMRank in a significant number of cases for short documents. These results suggest that the simplicity of our hypothesis vector is more effective when using smaller models, while larger models may require more sophisticated vectors to achieve optimal performance.
Appendix C Results of Non-Parametric SAMRank
@@
We implemented the non-parametric SAMRank method using the model LLAMA 3-8B, evaluating their three different scores to ensure that the non-parametric version benefits from these scores similarly to the original SAMRank. Table  7 shows the performance achieved by the non-parametric SAMRank in the three cases.
@@
Appendix D Most Relevant SAMs identified by Attention-Seeker
@@
Figure  4 shows the distribution of $R^{lh}$ scores for each layer, with the scores normalized for consistent comparison. For better visualization, the top 5% of outliers are filtered out from each distribution, distributions are color-coded according to their relevance rank, and a red line is added at $Relevance=3.125$ to indicate a uniform distribution where all layers are equally relevant. The analysis shows that the heads between layers 8 and 18, especially layers 9, 12, 14, and 15, received the highest relevance scores. This result suggests that heads located in intermediate layers of an LLM tend to specialize in processing tasks related to noun-noun relationships and selecting the most relevant nouns/phrases.
@@
Figure  5 supports the previous observation by illustrating the relevance scores $R^{lh}$ assigned to each head for a sample from the Inspec dataset. The score distribution across layers aligns with the general pattern in Figure  4, where most heads in the intermediate layers contribute significantly to keyphrase extraction, while only a few heads in the first and last layers exhibit high relevance. For the same Inspec sample, we compared the attention scores of candidate phrase tokens in layer 3 (lowest overall relevance score) and layer 14 (highest overall relevance score). Figure  6 shows the difference between these two layers: Layer 14 assigns higher attention scores to all candidates, with a clearer contrast between phrases, while layer 3 tends to focus on non-candidate tokens, resulting in lower, more uniformly distributed attention scores.
Model Method Inspec SemEval2017 SemEval2010 Krapivin
Small-sized Language Models
BERT SAMRank 0.04 0.06 6.76 6.74
Attention-Rank 0.12 0.13 7.89 7.89
GPT-2 SAMRank 0.05 0.06 7.16 7.33
Attention-Rank 0.12 0.13 7.90 8.11
T5 PromptRank 2.25 2.51 3.44 8.85
Medium-sized Language Models
LLAMA 3-8B SAMRank 0.22 0.30 11.98 12.05
Attention-Rank 0.68 0.76 14.00 14.01
PromptRank 5.88 9.89 45.12 42.82
Table 8: Average computation time (seconds) required by SAMRank, Attention-Rank, and PromptRank on small and medium-sized Language Models to process one document (batch size = 1).
Appendix E Computational cost
We evaluated the computation time required for the three methods with the best general performance in unsupervised keyphrase extraction: SAMRank, PromptRank, and Attention-Seeker. The computations were performed on an NVIDIA RTX A6000 GPU with 48 GB of memory. In all implementations, documents were processed individually (batch size = 1). We conducted the evaluation using four Language Models: BERT, GPT-2, and T5 for small-sized models, and LLAMA 3-8B for medium-sized models.
The results are shown in Table 8. PromptRank generally requires more computation time than the other two methods. This is due to its specific processing requirements: it extracts the logit for each "document" + "candidate" combination, effectively processing almost the same text $n$ times, where $n$ is the number of keyphrase candidates. A potential solution is to process all combinations as a batch, but this would shift the computational burden from time to memory resources.
Among the attention-based methods, SAMRank shows better computational efficiency compared to Attention-Seeker. This difference is expected because Attention-Seeker builds on SAMRank by introducing additional computations, including vector similarity calculations and linear transformations of SAMs. Despite these additional steps, both methods have comparable memory consumption, are efficient enough for real-time processing of short documents, and show similar computation times for long documents. Due to its superior performance, Attention-Seeker remains the preferred option overall. However, in scenarios where processing speed is critical for short documents and a small trade-off in F-score (about 2%) is acceptable, the non-parametric SAMRank would be the better choice.